OpenStudy (anonymous):

I need help finding the inverse of f(x)=3-2(x-1)^3 I need help finding the inverse of f(x)=3-2(x-1)^3 @Mathematics

OpenStudy (anonymous):

write f(x) as y and then swap the x and the y y = 3-2(x-1)^3 x = 3-2(y-1)^3 now solve for y again. \[y = \sqrt[3]{(x-3)/-2}+1\]

OpenStudy (anonymous):

ty using http://www.mathway.com/ it helps me

OpenStudy (anonymous):

you should probably write your final answer as \[y=\sqrt[3]{(3-x)/2}+1\] just to get rid of the negative and make it pretty.

OpenStudy (anonymous):

*try

OpenStudy (anonymous):

put \[x=3-2(y-1)^3\] and solve for y via \[2(y-1)^3=3-x\] \[(y-1)^3=\frac{3-x}{2}\] \[y-1=\sqrt[3]{\frac{3-x}{2}}\] \[y=\sqrt[3]{\frac{3-x}{2}}+1\] and so \[f^{-1}(x)=\sqrt[3]{\frac{3-x}{2}}+1\]

OpenStudy (anonymous):

what godspeed said